OpenStudy (anonymous):

What is the value for x in the proportion x over x plus ten equals seven over twelve?

OpenStudy (anonymous):

\[\frac{ x }{ x+10 }=\frac{ 7 }{ 12 }\]

OpenStudy (anonymous):

cross multiply to get 12x = 7x + 70 now solve for x

OpenStudy (anonymous):



OpenStudy (anonymous):

so we cross multiply 12x=7(x+10) distribute 7 across the expressino 12x=7x+70 substract 7x from both sides 5x=70 divide by 5 x=14

OpenStudy (anonymous):

you can teach or you can do homework for people